# Holoported Characters: Real-time Free-viewpoint Rendering of Humans from Sparse RGB Cameras

### v1

@@ Abstract
We present the first approach to render highly realistic free-viewpoint videos of a human actor in general apparel, from sparse multi-view recording to display, in real-time at an unprecedented 4K resolution. At inference, our method only requires four camera views of the moving actor and the respective 3D skeletal pose. It handles actors in wide clothing, and reproduces even fine-scale dynamic detail, e.g. clothing wrinkles, face expressions, and hand gestures. At training time, our learning-based approach expects dense multi-view video and a rigged static surface scan of the actor. Our method comprises three main stages. Stage 1 is a skeleton-driven neural approach for high-quality capture of the detailed dynamic mesh geometry. Stage 2 is a novel solution to create a view-dependent texture using four test-time camera views as input. Finally, stage 3 comprises a new image-based refinement network rendering the final 4K image given the output from the previous stages. Our approach establishes a new benchmark for real-time rendering resolution and quality using sparse input camera views, unlocking possibilities for immersive telepresence. More results can be found here: https://vcai.mpi-inf.mpg.de/projects/holochar/.
@@ 1 Introduction
Human free-viewpoint rendering is a long-standing and highly challenging problem in Vision and Graphics. The goal is to render any virtual viewpoint of the character given a discrete set of input camera views. Earlier approaches resorted to variants of explicit multi-view photometric reconstruction [38, 39, 5], reconstruction based on light fields [25], point primitives [45], or template-based scene representations [48, 2, 6] to compute an estimate of dynamic shape and appearance from the multi-view input. Despite great progress, these solutions are often limited along multiple dimensions, for example: They often require a very high number of camera views for good quality; computation times are often far from real-time; ghosting artifacts in rendered appearance are prevalent, since even the best methods fail to capture error-free scene geometry.
@@ 2 Related Work
Methods for novel view synthesis of general non-rigid scenes [37, 19, 24, 40] can be applied to videos with humans but struggle with large articulations due to the absence of human-specific priors. They enable scene replay but do not allow changes in the human pose without model retraining. Other human-specific replay methods can be trained using monocular [46] or multi-view videos [11, 54]. The visual quality of monocular methods can suffer due to the lack of explicit 3D information [46]. Multi-view methods [11, 54] can also render complex appearance elements like hair and clothing, even in real-time [54]. However, as they can solely replay the multi-view sequence they were trained on, they are not well suited for interactive telepresence applications, which is the focus of this work.
@@
Recent approaches generate novel 2D views of humans in novel poses but do not allow freely changing the 3D viewpoint [3, 34, 13]. Specialized approaches for the 3D free-viewpoint rendering of humans from monocular or multi-view RGB videos achieve high-fidelity results [46, 49, 23, 12, 52]; some of them generalize to poses unseen during training [33, 8, 9, 30, 26, 18, 42]. Their major limitation is that details such as clothing wrinkles are blurred as the pose information is ambiguous: A single pose can induce a variety of wrinkles, which often happens during training and which causes averaging of fine appearance details. While HDHumans [9] can generate high-frequency wrinkles (due to the generative formulation), those are often hallucinated and are not consistent with the ground-truth observations. Moreover, many methods also strongly rely on a neural rendering component, which makes them slow and not suitable for immersive applications. DDC [8] is one of the few methods that run in real-time and enable animatable control over the character: It generates an explicit mesh and texture and supports loose clothing. However, it still suffers from blurred details as most other methods.
@@
Approaches that use images for novel view rendering of dynamic scenes could also be applied to our task [21, 50, 41]. The recent work ENeRF [21] is an interactive real-time approach for free-viewpoint rendering with a neural representation driven by a sparse set of multi-view images. However, due to the lack of human priors, the method suffers from multi-view-consistency artifacts and does not generalize well to new poses and views. In contrast, our method achieves view consistent and high-quality results in real-time as we explicitly account for the humanoid structure, i.e. its articulation and non-rigid deformations.
@@ 3.1 Deformable Character Model
The resultant deformations are applied to the template in canonical space, followed by posing using Dual Quaternion Skinning [14], yielding the final positioned vertex $\boldsymbol{v}_{i}\in\mathbb{R}^{3}$ for each vertex $i$ of the template mesh. The aggregate of these posed vertices $\boldsymbol{V}\in\mathbb{R}^{N\times 3}$ is derived by stacking each $\boldsymbol{v}_{i}$, where $N$ represents the number of vertices in the template. Their entire model is trained in multiple stages, using only multi-view data. Please refer to the supplementary material and [8] for more technical details.
@@
Additionally, we observe that the coarse deformation, which $f_{\mathrm{eg}}(M_{t})$ predicts, causes artifacts for the hands, as they are a part of the body with high articulation very close together, which cannot be modeled by coarse deformation. To address this issue, we set specific parameters at hand vertices to zero, preventing the model from deviating from the initial skinning deformation. This significantly improves the performance, as illustrated in Fig. 3.
@@ 4 Results
We evaluate our method on two sequences from the DynaCap dataset [8], including one subject wearing tight clothes and one subject in loose clothes. Since in DynaCap the actors are with fists clenched, we also recorded a new dataset with three novel sequences where the actors were instructed to use both hands in a natural manner; Therefore, we can evaluate the capability of our method to represent fingers. We captured our new sequences in a multi-view camera setup with 120 synced 4K resolution cameras. Each sequence is performed by an actor wearing different clothes and is split into around 20K frames for training and 7K frames for testing. The camera streams are further divided into training and testing views. To ensure a fair comparison to other methods and to avoid a bias in the results due motion tracking errors (which is not the focus of this work), we run motion capture on 20 cameras.
@@ 4.2 Comparison
Next, we compare to methods, that takes sparse multi-view images as input and render free-viewpoint videos in real time. ENeRF [21] is a method designed for general scenes, while DVA [31] is specifically designed for humans, thus, most closely related to our approach. Originally, ENeRF uses the nearest two views at test time from a dense setup to produce the target view, which effectively violates our sparse view requirement during inference. Instead, we retrain ENeRF using all the training views/frames and always perform inference using four fixed cameras, which is the same setting used for all methods. As shown in Fig. 7, ENeRF, because of their constrained color formulation, fails to maintain multi-view consistency and produces blurry results for novel views, especially for those that are far away from the fixed views. We further compare to DVA [31]. Since DVA’s character model is solely based on skinning and regularizes large deformations, it performs poorly in representing high-frequency details and loose types of apparel. We also found that their volume-primitive-based formulation is hard to train on a large variety of poses leading to a degradation in quality when training on large-scale data (such as the DynaCap dataset), which is essential to achieve desired pose generalization at test time. In consequence, we found that DVA produces blurred results and noticeable artifacts for loose types of apparel (see insets in Fig. 7). In contrast, our method results in high-quality renderings with sharp details, wrinkles, and it can also handle loose types of clothing. We refer to the supplemental material for a more detailed analysis of these methods.
@@
Method RT PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ (×$\times 1000$) FIDnormal-$\downarrow$ PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ (×$\times 1000$) FIDnormal-$\downarrow$
@@ 4.3 Ablation
Method PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ (×$\times 1000$) FIDnormal-$\downarrow$ Res.
@@
We also evaluate the effect of the Chamfer loss for the 3D surface supervision in Tab. 2. It significantly improves the rendering quality of our method, as it leads to surfaces that are better aligned with the input images.
@@
The super-resolution (SR) module significantly contributes to improving the visual quality and the metrics. Importantly, the SRNet removes observable noise potentially, which is due to the discrete mapping of texels, and the typical mesh boundary artifacts (see the inset “w/o SR” and “Ours” in Fig. 8). Considering the rendering resolution, we observe that training our approach on 4K images translates into better metrics for both evaluation cases (1K and 4K image resolution), especially when considering LPIPS and FID, which are more sensitive to high-frequency details in comparison to PSNR.
@@ 5 Conclusions
We introduced Holoported Characters, a novel method for real-time free-viewpoint rendering of humans from sparse RGB cameras and 3D skeletal poses. Our approach achieves unprecedented 4K resolution and runtime performance by seamlessly integrating neural and explicit components. It surpasses the previous state of the art, particularly excelling in challenging scenarios like loose clothing and diverse motions. We believe our work is an important step towards telepresence enabling immersive communication across the globe. For the future, we plan to explore multi-layered human representations potentially being able to model topological changes such as opening a jacket, which our method currently cannot handle.
@@ Appendix A Supplementary Material
We first present more technical details on our character model (Sec. A.1), and the depth testing approach we use for texel visibility (Sec. A.2). Then, we provide more implementation details (Sec. A.3) and additional results, and applications of our method (Sec. A.4, Sec. A.5). We further provide additional comparisons (Sec. A.6), ablations (Sec. A.7), and analysis of methods closest to our setting (Sec. A.8). Finally, we discuss the limitations of our approach (Sec. A.9).
@@ A.1 Deformable Character Model
Our character model takes in a temporal motion input ${M=\{(\boldsymbol{\theta}_{t-W},\boldsymbol{\alpha}_{t-W},\boldsymbol{z}_{t-W})....(\boldsymbol{\theta}_{t},\boldsymbol{\alpha}_{t},\boldsymbol{z}_{t})\}}$ and deforms a template mesh capable of modelling loose clothing. Here, $\boldsymbol{\theta}_{t}\in\mathbb{R}^{P},\boldsymbol{\alpha}_{t}\in\mathbb{R}^{3},\boldsymbol{z}_{t}\in\mathbb{R}^{3}$ refer to the skeleton DOFs, root translation, and root rotation, respectively. We leverage the explicit character representation of Habermann et al. [8], as it is differentiable, real-time, and can model loose clothing:
$C_{i}(\boldsymbol{\theta}_{t},\boldsymbol{\alpha}_{t},\boldsymbol{z}_{t},\boldsymbol{A},\boldsymbol{T},\boldsymbol{d}_{i})=\boldsymbol{v}_{i}.$ (13)
In their character formulation, the initial template is downsampled to an embedded graph $\boldsymbol{G}$ with $K$ nodes, and the parameters $\boldsymbol{A}\in\mathbb{R}^{K\times 3}$ and $\boldsymbol{T}\in\mathbb{R}^{K\times 3}$ are the rotation and translation of each of the $K$ nodes stacked on top of each other, describing the coarse deformation in canonical space. $\boldsymbol{d}_{i}\in\mathbb{R}^{3}$ is the per-vertex displacement in canonical space. The final location
@@
of a vertex in canonical space $\boldsymbol{y}_{i}\in\mathbb{R}^{3}$ is determined by the weighted addition of the rotation and translation of its neighbors in the embedded graph $N_{\mathrm{vn}(i)}$, and the per-vertex deformation $\boldsymbol{d}_{i}$ is added. Here, $w_{i,k}$ is the weight of the $i$th vertex assigned to node $k$, $R(\cdot)$ converts Euler angles to rotation matrices, $\boldsymbol{a}_{k}$ and $\boldsymbol{t}_{k}$ are the $k$th rows of $\boldsymbol{A}$ and $\boldsymbol{T}$, respectively, $\hat{\boldsymbol{v}}_{i}$ is the $i$th vertex of the template mesh, and $\boldsymbol{g}_{k}$ is the $k$th node of $\boldsymbol{G}$.
Their model predicts $\boldsymbol{A}$, $\boldsymbol{T}$, and $\boldsymbol{d}_{i}$ using structure-aware graph neural networks [8] $f_{\mathrm{eg}}(e(M))$ and $f_{\mathrm{delta}}(d(M))$, where $e(\cdot)$ and $d(\cdot)$ are their proposed motion to embedded graph embedding and motion to vertex embedding. Finally, they apply the skeleton pose to the deformed vertex in canonical space $\boldsymbol{y_{i}}$ to get the final deformed vertex location
@@
where the rotation $R_{\mathrm{sk},k}$, and translation $\boldsymbol{t}_{\mathrm{sk},k}$ for node $k$ are determined by Dual Quaternion skinning [15]. The final deformed vertices $\boldsymbol{V}\in\mathbb{R}^{N\times 3}$ are generated by stacking $\boldsymbol{v_{i}}$.
@@ A.2 Depth Testing
We solve the texel visibility problem by using a depth testing approach. We assign a depth to each texel ${T_{\mathrm{pos}}\in\mathbb{R}^{TW\times TH\times 3}}$, using the barycentric coordinates of the face it is projected onto. Then, we compute the depth ${D_{c,V}\in\mathbb{R}^{H\times W\times 3}}$ and the texel to pixel mapping ${F_{\mathrm{warp_{c,V}}}:(u,v)\rightarrow(x,y)}$ for a view $c$ using differentiable rasterisation from [7], camera parameters, and the deformed vertex positions $\boldsymbol{V}$. If the depth at a texel $(u,v)$ is within a $\epsilon$ norm ball to the depth of the pixel it lands on, we mark it as visible:
@@
A.3 Implementation Details
The obtained multi-view frames are processed with background-foreground segmentation [22] and per-frame mesh reconstruction using NeuS2 [43]. Motion tracking is obtained with a commercial markerless capture system [36]. Our method is implemented in TensorFlow [1] and trained using the Adam optimizer [16] with a constant learning rate of $10^{-4}$ until convergence.
The character model and the TexFeatNet are supervised on images with a resolution of $1028\times{752}$ pixels, and the SRNet is trained with a full image resolution of $4112\times{3008}$ pixels. The projective texturing module uses images of resolution $4112\times{3008}$ as input, and the texture maps are generated at a resolution of $1024\times{1024}$ pixels. For the Chamfer loss, we randomly sample 10K points from the reconstructed surface for every frame. TexFeatNet utilizes a UNet architecture [32], and SRNet is implemented as a shallow architecture with enhanced residual blocks [20]. Our complete framework is trained using two Nvidia A100 GPUs with 80GB memory and a batch size of four. Tab. 3 provides a component-wise runtime breakdown.
Component FPSnormal-$\uparrow$ Latencynormal-$\downarrow$
Character Model 100 0.010s
Projective Texturing 27 0.037s
TexfeatNet 100 0.010s
SRNet 31.25 0.032s
Table 3: Runtime Breakdown. Here we present a component wise throughput breakdown of our pipeline. Note that all of our components run within the real-time limit of $25$ frames per second (FPS).
Figure 9: Qualitative Ablations. Note that our design choices for the SRNet module, lead to better results qualitatively, especially at the borders of the human.
Figure 10: Application: Texture Editing. As we use explicit textures as an underlying latent for appearance, and they are temporally spatially aligned, we can perform edits such as adding a logo to a character.
Method PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ (×$\times 1000$) FIDnormal-$\downarrow$ Res.

### v2

@@ Abstract
We present the first approach to render highly realistic free-viewpoint videos of a human actor in general apparel, from sparse multi-view recording to display, in real-time at an unprecedented 4K resolution. At inference, our method only requires four camera views of the moving actor and the respective 3D skeletal pose. It handles actors in wide clothing, and reproduces even fine-scale dynamic detail, e.g. clothing wrinkles, face expressions, and hand gestures. At training time, our learning-based approach expects dense multi-view video and a rigged static surface scan of the actor. Our method comprises three main stages. Stage 1 is a skeleton-driven neural approach for high-quality capture of the detailed dynamic mesh geometry. Stage 2 is a novel solution to create a view-dependent texture using four test-time camera views as input. Finally, stage 3 comprises a new image-based refinement network rendering the final 4K image given the output from the previous stages. Our approach establishes a new benchmark for real-time rendering resolution and quality using sparse input camera views, unlocking possibilities for immersive telepresence. Code and data is available on our project page.
@@ 1 Introduction
Human free-viewpoint rendering is a long-standing and highly challenging problem in Vision and Graphics. The goal is to render any virtual viewpoint of the character given a discrete set of input camera views. Earlier approaches resorted to variants of explicit multi-view photometric reconstruction [45, 46, 5], reconstruction based on light fields [29], point primitives [52], or template-based scene representations [55, 2, 6] to compute an estimate of dynamic shape and appearance from the multi-view input. Despite great progress, these solutions are often limited along multiple dimensions, for example: They often require a very high number of camera views for good quality; computation times are often far from real-time; ghosting artifacts in rendered appearance are prevalent since even the best methods fail to capture error-free scene geometry.
@@ 2 Related Work
Methods for novel view synthesis of general non-rigid scenes [44, 21, 28, 47] can be applied to videos with humans but struggle with large articulations due to the absence of human-specific priors. They enable scene replay but do not allow changes in the human pose without model retraining. Other human-specific replay methods can be trained using monocular [53] or multi-view videos [12, 61]. The visual quality of monocular methods can suffer due to the lack of explicit 3D information [53]. Multi-view methods [12, 61] can also render complex appearance elements like hair and clothing, even in real-time [61]. However, as they can only replay the multi-view sequence they were trained on, they are not well suited for interactive telepresence applications, which is the focus of this work.
@@
Recent approaches generate novel 2D views of humans in novel poses but do not allow freely changing the 3D viewpoint [3, 39, 14, 25, 26]. Specialized approaches for the 3D free-viewpoint rendering of humans from monocular or multi-view RGB videos achieve high-fidelity results [53, 56, 27, 13, 59]; some of them generalize to poses unseen during training [38, 8, 10, 35, 30, 20, 49]. Their major limitation is that details such as clothing wrinkles are blurred as the pose information is ambiguous: A single pose can induce a variety of wrinkles, which often happens during training and which causes averaging of fine appearance details. While HDHumans [10] and other concurrent works [19, 31] can generate high-frequency wrinkles (due to the generative formulation), those are often hallucinated and are not consistent with the ground-truth observations. Moreover, many methods also strongly rely on a neural rendering component, which makes them slow and not suitable for immersive applications. DDC [8] is one of the few methods which run in real-time and enable animatable control over the character: It generates an explicit mesh and texture and supports loose clothing. However, it still suffers from blurred details as most other methods.
@@
Approaches that use images for novel view rendering of dynamic scenes could also be applied to our task [23, 57, 48]. The recent work ENeRF [23] is an interactive and real-time approach for free-viewpoint rendering with a neural representation driven by a sparse set of multi-view images. However, due to the lack of human priors, the method suffers from multi-view-consistency artifacts and does not generalize well to new poses and views. In contrast, our method achieves view consistent and high-quality results in real-time as we explicitly account for the humanoid structure, i.e. its articulation and non-rigid deformations.
@@ 3.1 Deformable Character Model
The resultant deformations are applied to the template in canonical space, followed by posing using Dual Quaternion Skinning [15], yielding the final positioned vertex $\boldsymbol{v}_{i}\in\mathbb{R}^{3}$ for each vertex $i$ of the template mesh. The aggregate of these posed vertices $\boldsymbol{V}\in\mathbb{R}^{N\times 3}$ is derived by stacking each $\boldsymbol{v}_{i}$, where $N$ represents the number of vertices in the template. Their entire model is trained in multiple stages, using only multi-view data. Please refer to the supplementary material and the original work [8] for more technical details.
@@
Additionally, we observe that the coarse deformation, which $f_{\mathrm{eg}}(M_{t})$ predicts, causes artifacts for the hands, as they are a part of the body with high articulation very close together, which cannot be modeled by coarse deformation. To address this issue, we set specific parameters at hand vertices to zero, preventing the model to deviate from the initial skinning deformation. This significantly improves the performance, as illustrated in Fig. 3.
@@ 4 Results
We evaluate our method on two sequences from the DynaCap dataset [8], including one subject wearing tight clothes and one subject in loose clothes. Since in DynaCap the actors are with fists clenched, we also recorded a new dataset with three novel sequences where the actors were instructed to use both hands in a natural manner; therefore, we can evaluate the capability of our method to represent fingers. We captured our new sequences in a multi-view camera setup with 120 synced 4K resolution cameras. Each sequence is performed by an actor wearing different clothes and is split into around 20K frames for training and 7K frames for testing. The camera streams are further divided into training and testing views. To ensure a fair comparison to other methods and to avoid a bias in the results due motion tracking errors (which is not the focus of this work), we run markerless motion capture on 20 cameras.
@@ 4.2 Comparison
Next, we compare to methods, that take sparse multi-view images as input and render free-viewpoint videos in real time. ENeRF [23] is a method designed for general scenes, while DVA [36] is specifically designed for humans, thus, most closely related to our approach. Originally, ENeRF uses the nearest two views at test time from a dense setup to produce the target view, which effectively violates our sparse view requirement during inference. Instead, we retrain ENeRF using all the training views/frames and always perform inference using four fixed cameras, which is the same setting used for all methods. As shown in Fig. 7, ENeRF, because of their constrained color formulation, fails to maintain multi-view consistency and produces blurry results for novel views, especially for those that are far away from the input views. We further compare to DVA [36]. Since DVA’s character model is solely based on skinning and regularizes large deformations, it performs poorly in representing high-frequency details and loose types of apparel. We also found that their volume-primitive-based formulation is hard to train on a large variety of poses leading to a degradation in quality when training on large-scale data (such as the DynaCap dataset), which is essential to achieve desired pose generalization at test time. In consequence, we found that DVA produces blurred results and noticeable artifacts (see insets in Fig. 7). In contrast, our method results in high-quality renderings with sharp details, wrinkles, and it can also handle loose clothing. We refer to the supplemental for a more detailed analysis of these methods.
@@
[…]
@@ 4.3 Ablation
[…]
@@
We also evaluate the effect of the Chamfer loss for the 3D surface supervision in Tab. 2. It significantly improves the rendering quality, as it leads to surfaces that are better aligned with the input images.
@@
The super-resolution (SR) module significantly contributes to improving the visual quality and the metrics. Importantly, the SRNet removes observable noise, which is due to the discrete mapping of texels, and the typical mesh boundary artifacts (see the inset “w/o SR” and “Ours” in Fig. 8). Considering the rendering resolution, we observe that training our approach on 4K images translates into better metrics for both evaluation cases (1K and 4K image resolution), especially when considering LPIPS and FID, which are more sensitive to high-frequency details in comparison to PSNR.
@@ 5 Conclusions
We introduced Holoported Characters, a novel method for real-time free-viewpoint rendering of humans from sparse RGB cameras and 3D skeletal poses. Our approach achieves unprecedented 4K resolution and runtime performance by seamlessly integrating neural and explicit components. We believe our work is an important step towards telepresence enabling immersive communication across the globe. However, our method is not without limitations. For example, we cannot handle topological changes such as opening a jacket. For the future, we plan to explore multi-layered human representations potentially being able to model topological changes, which our method currently cannot handle.
Acknowledgments.
This project was supported by the Saarbrücken Research Center for Visual Computing, Interaction and AI. Christian Theobalt was supported by ERC Consolidator Grant 4DReply (770784).
@@ Appendix A Supplementary Material
We first present more technical details on our character model (Sec. A.1), and the depth testing approach for texel visibility (Sec. A.2). Then, we provide more implementation as well as data processing details (Sec. A.4 and Sec. A.3), additional results, and applications of our method (Sec. A.5, Sec. A.6). We further show additional comparisons (Sec. B), ablations (Sec. B.1), and analysis of methods closest to our setting (Sec. B.2). Finally, we talk about the limitations of our approach (Sec. B.3).
@@ A.1 Deformable Character Model
Our character model takes a temporal motion $M=\{(\boldsymbol{\theta}_{t-W},\boldsymbol{\alpha}_{t-W},\boldsymbol{z}_{t-W})....(\boldsymbol{\theta}_{t},\boldsymbol{\alpha}_{t},\boldsymbol{z}_{t})\}$ as input and deforms a template mesh capable of modelling loose clothing. Here $\boldsymbol{\theta}_{t}\in\mathbb{R}^{P},\boldsymbol{\alpha}_{t}\in\mathbb{R}^{3},\boldsymbol{z}_{t}\in\mathbb{R}^{3}$ refer to the skeleton degrees of freedom, root translation, and root rotation, respectively. We leverage the explicit character representation of Habermann et al. [8]
$C_{i}(\boldsymbol{\theta}_{t},\boldsymbol{\alpha}_{t},\boldsymbol{z}_{t},\boldsymbol{A},\boldsymbol{T},\boldsymbol{d}_{i})=\boldsymbol{v}_{i},$ (13)
as it is differentiable, real-time, and models loose clothing. In their character formulation, the initial template $T$ is downsampled to an embedded graph [41] $G$ with $K$ nodes, and the parameters $\boldsymbol{A}\in\mathbb{R}^{K\times 3},\boldsymbol{T}\in\mathbb{R}^{K\times 3}$ are the rotation and translation of each of the $K$ nodes stacked on top of each other, describing the coarse deformation of $T$ in canonical space. $\boldsymbol{d}_{i}\in\mathbb{R}^{3}$ is the per-vertex displacement in canonical space. The final location
@@
of a vertex in canonical space $\boldsymbol{y}_{i}\in\mathbb{R}^{3}$ is determined by the weighted addition of the rotation and translation of its neighbours in the embedded graph $N_{\mathrm{vn}(i)}$, and finally adding the per-vertex deformation $\boldsymbol{d}_{i}$. Here, $w_{i,k}$ is the weight the $i$th vertex assigns to node k. $R(\cdot)$ is the function that converts Euler angles to matrices. $\boldsymbol{a}_{k}$, $\boldsymbol{t}_{k}$ are $k$th rows of $\boldsymbol{A}$, $\boldsymbol{T}$ respectively. $\hat{\boldsymbol{v}}_{i}$ is the $i$th vertex of $T$ and $\boldsymbol{g}_{k}$ is the $k$th node of G.
Their model predicts $\boldsymbol{A}$, $\boldsymbol{T}$, and $\boldsymbol{d}_{i}$ using structure-aware graph neural networks[8], referred to as $f_{\mathrm{eg}}(e(M))$ and $f_{\mathrm{delta}}(d(M))$. $e(\cdot)$ and $d(\cdot)$ are their proposed motion to embedded graph embedding and motion to vertex embedding. Finally, they apply the skeleton pose to the deformed vertex in canonical space $\boldsymbol{y_{i}}$ to obtain the final deformed vertex location
@@
where the rotation $R_{\mathrm{sk},k}$, and translation $\boldsymbol{t}_{\mathrm{sk},k}$ for node $k$ are determined by Dual Quaternion Skinning [16]. The vertex matrix $\boldsymbol{V}\in\mathbb{R}^{N\times 3}$ can be obtained by stacking $\boldsymbol{v_{i}}$.
@@ A.2 Depth Testing
We address the texel visibility problem by using a depth testing approach. We assign a 3D depth to each texel $T_{\mathrm{pos}}\in\mathbb{R}^{TW\times TH\times 3}$, using the barycentric coordinates of the face it lands on. Then, we compute the depth $D_{c,V}\in\mathbb{R}^{H\times W\times 3}$ and the texel to pixel mapping $F_{\mathrm{warp_{c,V}}}:(u,v)\rightarrow(x,y)$ for a particular view $c$ using differentiable rasterisation from DeepCap [7, 9] given the camera parameters and the deformed vertex positions $\boldsymbol{V}$. If the depth at a particular texel $(u,v)$ is within a $\epsilon$ norm ball to the depth of the pixel it lands on, we mark it as visible, i.e.
@@
Novel View Novel Pose
[…]
DDC 11.53 11.24 15.5 15.6
HDHumans 11.52 11.21 13.7 13.5
Ours 10.4 9.3 13.1 12.9
Table 3: Quantitative Comparison on Surface Distance. We evaluate the surface tracking in the novel view and novel pose setting on subject S1 and report the Chamfer (CD) and Hausdorff (HD) distance with respect to the ground truth. Note that our approach outperforms previous works, thanks to the additional Chamfer penalty. This is significant in enhancing the quality of our projective texturing pipeline. Error reported in $mm$.
[…]
Table 4: Runtime Breakdown. Here, we present a component-wise runtime breakdown of our pipeline in terms of frames per second (fps) and latency (in seconds $s$). Note that all of our components run within the real-time limit of 25 fps.
A.3 Additional Data Processing Details
Each actor in our dataset is scanned using a commercially available 3D scanner [43] where the mesh is obtained from multi-view stereo reconstruction [^1] https://www.agisoft.com. Following this, we downsample the high-resolution mesh to around $9000$ faces for each character. For character rigging, we apply markerless motion capture [42] on the multi-view images from the scanner to obtain the skeletal pose. Given the pose and the template scan, we apply Blender’s [^2] https://www.blender.org automated skinning weight computation. The UV parameterization is obtained from a photometric stereo reconstruction software. However, the effect of UV parameterization and optimizing the UV parameterization is a future work that we believe merits further investigation.
A.4 Implementation Details
The obtained multi-view frames are processed with foreground segmentation [24] and per-frame mesh reconstruction using NeuS2 [50]. Motion tracking is obtained with a commercial markerless capture system [42]. Our method is implemented in TensorFlow [1] and trained using the Adam optimizer [17] with a constant learning rate of $10^{-4}$ until convergence.
The character model and the TexFeatNet are supervised on images with a resolution of $1028\times{752}$ pixels, and the SRNet is trained with a full image resolution of $4112\times{3008}$ pixels. The projective texturing module uses images of resolution $4112\times{3008}$ as input, and the texture maps are generated at a resolution of $1024\times{1024}$ pixels. We randomly sample 10K points from the reconstructed surface for the Chamfer loss for every frame.
TexFeatNet utilizes a UNet architecture [37], and SRNet is implemented as a shallow architecture with enhanced residual blocks [22]. Our complete framework is trained using two Nvidia A100 GPUs with 80GB memory and a batch size of four. Tab. 4 provides a runtime breakdown of our modules.
Figure 9: Qualitative Ablations. Note that our design choices for the SRNet module, lead to qualitatively better results, especially at the borders of the human.
Figure 10: Application: Texture Editing. As we use explicit textures as an underlying latent for appearance, and they are temporally and spatially aligned, we can perform 2D texture edits such as adding a logo onto the character’s shirt.
[…]